# Unsupervised Visual Chain-of-Thought Reasoning via Preference Optimization

### v1

@@ 4 Experiments
In Sec. 4.1, we outline our experimental setup, including datasets, evaluation metrics, baselines, and implementation details. Sec. 4.2 evaluates UV-CoT on six widely used datasets, demonstrating its superiority over existing baselines. Sec. 4.4 presents ablation studies analyzing impact of key components on overall performance. In Sec. 4.3, we assess the generalization ability of UV-CoT in zero-shot settings across four diverse datasets. In Sec. 4.5, we evaluate the quality of generated bounding boxes across all datasets. In Sec. 4.6, we analyze how our evaluation process reduces task difficulty. In Sec. 4.7, we provide additional analyses on token efficiency, visualization.
4.1 Experimental Setup
Datasets. To generate diverse image-level CoT data, we select nine datasets spanning five domains:
  • Text/Document: DocVQA [21], TextVQA [29], DUDE [33], and SROIE [10].
  • Chart: InfographicsVQA [22].
  • General VQA: Flickr30k [25] and Visual7W [44].
  • Relation Reasoning: VSR [16] and GQA [11].
  • High-Resolution Image Reasoning: V Bench [35].
We follow the procedure in Sec. 3.1 to generate high-quality preference data based on these datasets. Notably, we also provide a model trained on data excluding DUDE, SROIE, and Visual7W, which is used to evaluate zero-shot performance. For details on dataset composition and preprocessing, please refer to Appendix C.1.
Evaluation. Following the evaluation methodology of prior work [18], we use GPT-4o [23] as the evaluator. GPT-4o is prompted to assign a numerical score between 0 and 1, with higher scores indicating better prediction accuracy. Details on prompt design and setup are provided in Appendix C.2.
Baselines We compare our model against five state-of-the-art baselines. Specifically, LLaVA-1.5-7B [17], LLaVA-1.5-13B and OmniLMM-12B serve as strong comprehensive baselines, as they are pretrained on large-scale multimodal data and fine-tuned with high-quality instruction data. OmniLMM-12B is also our evaluator model. Additionally, MiniCPM-o-8B [38] employs an adaptive visual encoding mechanism to capture fine-grained features, improving visual understanding. Finally, Visual-CoT-7B [28] utilizes labeled data to learn image-level CoT via SFT, achieving competitive performance.
Implementation Details. We use LLaVA-1.5-7B as the target model and OmniLMM-12B as the evaluator. To ensure scalability, we avoid using GPT-4 as the evaluator, preventing constraints imposed by high API costs. Our UV-CoT training follows a two-stage process aligned with the LLaVA-1.5 training framework. In the first stage, the vision encoder and language model weights are frozen while training on image-text caption data. In the second stage, all weights are unfrozen for full training. We implement iterative learning of UV-CoT over four iterations, utilizing a total of 249K preference data pairs. Notably, UV-CoT achieves higher data efficiency than Visual-CoT, which uses 376K data pairs. For each iteration, we train the model with AdamW optimizer for 4 epochs with a learning rate of $0.15\times 10^{-7},\beta=$ 0.1, and a batch size of 8. In total, data generation takes 80 hours and training requires 60 hours, both conducted on an 8$\times$×A100 40GB machine. Additionally, we provide a variant UV-CoT trained with extra SFT on 10% of the labeled Visual-CoT data, denoted as UV-CoT (10%).
4.2 Comparisons to State-of-The-Arts
@@
  • Both image-level CoT models, Visual-CoT and our UV-CoT, achieve a significant performance gain over all baselines, even outperforming the larger LLaVA-1.5-13B model across all datasets. This highlights the effectiveness of image-level CoT reasoning. Although MiniCPM-o-8B employs crop-and-select strategies to enhance focus on specific image regions, this approach remains limited. In contrast, image-level CoT fully utilizes MLLMs to generate key regions, whereas MiniCPM-o-8B relies on rule-based cropping applied to only a few layers. Nonetheless, MiniCPM-o-8B outperforms LLaVA-1.5-7B by an average of 3.5%, underscoring the importance of explicitly integrating visual cues.
  • UV-CoT significantly outperforms OmniLMM-12B by an average of 5.1% across all evaluated datasets. This underscores that our method extend beyond merely distilling knowledge from larger evaluator model. Directly generating accurate bounding boxes is a challenging task for MLLMs, requiring precise spatial understanding and localization. UV-CoT addresses this challenge by reformulating the task to focus on evaluating the final answer, which is inherently more tractable.
@@
  • UV-CoT outperforms Visual-CoT-7B on TextVQA (+1.3%) and VSR (+1.6%). Moreover, UV-CoT (10%) surpasses Visual-CoT-7B across four datasets and 2.1% in average, with particularly notable improvements on TextVQA (+2.5%), GQA (+2.2%), and VSR (+2.1%). On the remaining two datasets, UV-CoT (10%) achieves comparable performance to Visual-CoT-7B. Notably, our method is trained on significantly less unlabeled data (249K) than Visual-CoT-7B, which uses 376K labeled CoT data. This success can be attributed to our high-quality preference data generation and the enhanced preference optimization method, which further improves image-level CoT learning.
4.3 Zero-Shot Performance (Generalization)
@@
Comprehensive Zero-Shot. We evaluate our model’s zero-shot image-level CoT capabilities using the test splits of SROIE, DUDE, and Visual7W, without training on these datasets. Since our method does not require labeled data, we can effectively evaluate these datasets even in the absence of bounding box annotations. Specifically, we train a variant of our model, denoted as UV-CoT, using preference data that incorporates these datasets. The results are presented in Tab. 2. UV-CoT consistently outperforms all baselines across all zero-shot datasets (2.5% in average), achieving notable improvements on DUDE (+3.5%) and Visual7W (+3.5%), which underscores its strong generalization ability. In contrast, supervised CoT learning, which depends on labeled data, often overfits to specific annotation distributions, limiting its ability to generalize to new tasks. UV-CoT, however, leverages preference optimization by focusing on relative comparisons rather than absolute labels, thereby reducing reliance on specific annotated distributions. Furthermore, UV-CoT achieves even greater improvements (5.1% in average), with a particularly significant margin on SROIE (+4.6%), demonstrating that our model can effectively learn the image-level CoT process without depending on labeled data. Based on this capability, our method can be easily extended to additional tasks, eliminating the need for costly human annotation.
High-Resolution Image Reasoning. To evaluate the capability in handling challenging high-resolution image reasoning tasks, we test UV-CoT on the V Bench [35], which comprises 238 images from the SA-1B dataset [12] with an average resolution of 2246$\times$1582. We conduct experiments on three representative tasks—Attributes (object attribute recognition), GPT4V-Hard (complex visual reasoning), and OCR (text recognition)-with the results summarized in Tab. 3. CoT-based methods, UV-CoT (+19.8% on average) and Visual-CoT-7B (+14.3% on average), significantly outperform non-CoT baselines, especially for OCR tasks (more than 50%). And UV-CoT achieves the highest performance across all tasks, surpassing Visual-CoT-7B by 5.5% on average. This illustrates the effectiveness of the visual CoT in high-resolution image reasoning. The superior performance over Visual-CoT-7B further demonstrates that UV-CoT optimizes the visual CoT learning process, enabling fine-grained region-based reasoning.
@@ 4.4 Ablation Studies
This section presents ablation experiments analyzing key components, as shown in Tab. 4.
Image-Level CoT: To evaluate the effectiveness of image-level CoT reasoning, we introduce two variants. w/o UV-CoT removes the CoT reasoning process, relying on direct inference without intermediate reasoning steps. UV-CoT w/ G.T. BBox utilizes ground-truth (GT) bounding boxes in the first step of the CoT process to assess the impact of accurate region selection. Results show that w/o UV-CoT leads to a significant performance drop across all datasets (7.7% in average), demonstrating the effectiveness of our image-level CoT approach. Meanwhile, UV-CoT w/ G.T. BBox serves as an upper bound for our method. In Flickr30k, our model performs on par with UV-CoT w/ G.T. BBox, indicating that it accurately selects key regions. However, in DocVQA and Infographics VQA, UV-CoT w/ G.T. BBox significantly outperforms our method, highlighting the challenge of generating precise bounding boxes. This suggests that our approach has a high upper limit and strong potential. Future work can focus on improving bounding box accuracy, especially for complex tasks.
Preference Optimization: w/ naive DPO applies a standard DPO loss. Its performance deteriorates across all datasets (1.9% in average), with particularly significant drops on DocVQA (2.5%) and Flickr30k (2.6%), highlighting the limitations of standard DPO in image-level CoT learning. This result highlights the advantages of sDPO, which enhances preference differentiation by incorporating scores. The consistent performance gains further validate the accuracy of our preference scores.
Iterative Learning: w/o iterative learning generates preference pairs for the entire set of question queries $\mathcal{X}$ in a single pass and trains once, leading to a significant performance drop (3.5% in average) across evaluated datasets. This underscores the importance of iterative learning in dynamically adapting the preference data distribution to remain aligned with the model’s evolving policy throughout the training process.
Response Evaluation: We assess the impact of the response evaluation stage by setting $\gamma=0$, resulting in the w/o $\gamma$ model. In this setting, the generated preference scores ignore the influence of the next response. This leads to a significant performance drop across all datasets (8.8% in average), particularly on TextVQA (17.2%), highlighting the challenge that MLLMs face in directly evaluating bounding boxes. This result underscores the necessity of incorporating next response in our evaluation method.
Figure 3: Object detection performance comparisons on (a) training datasets and (b) zero-shot datasets.
Model DocVQA TextVQA InfographicsVQA Flickr30k GQA VSR Average OmniLMM-12B 0.254 0.578 0.172 0.621 0.509 0.523 0.443 IF: OmniLMM-12B + CoT 0.305 0.722 0.217 0.675 0.580 0.595 0.516 LLaVA-1.5-7B 0.198 0.507 0.131 0.539 0.480 0.504 0.393 IF: LLaVA-1.5-7B + CoT 0.245 0.577 0.149 0.606 0.530 0.533 0.440 Tr: UV-CoT (Evaluator: self-evaluated) 0.242 0.609 0.153 0.598 0.517 0.526 0.441 Tr: UV-CoT (Evaluator: OmniLMM-12B) 0.265 0.686 0.173 0.632 0.536 0.548 0.473
Table 5: Analysis of evaluator model. + CoT refers to inference assisted by bounding boxes generated by UV-CoT. Self-evaluated denotes using target model as the evaluator during training. IF and Tr indicate experiments conducted on the inference and training process, respectively.
Figure 4: The model performance comparison regarding different sizes of visual tokens.
Figure 5: Visualization of preference data generated by Algorithm 1. Preferred bounding boxes are shown in red. Dis-preferred bounding boxes are in blue.
@@
4.5 Object Detection Performance
Previous experiments have demonstrated the exceptional performance of our image-level CoT process, highlighting the effectiveness of the generated bounding boxes in improving subsequent answers. However, we remain interested in evaluating the quality of the bounding boxes themselves. To this end, we employ GPT-4o to score the model-generated bounding boxes on both training datasets (Fig. 3 (a))and zero-shot datasets (Fig. 3 (b)), where best performances are marked with stars. InforVQA is a short for InfographicsVQA. We have the following observations:
  • Our UV-CoT outperforms Visual-CoT-7B, achieving higher scores in five out of six training datasets, supporting its superior performance in generating helpful bounding box.
  • The bounding box quality is closely related to the final performance. Our model exhibits lower scores (below 0.210) for bounding box generation in DocVQA and InfographicsVQA, correlating with its reduced final scores in these datasets ((below 0.290 in Tab. 1). It underscores the validity of evaluating bounding box quality through its impact on subsequent answers. In addition, this observation explains the performance gap between UV-CoT and UV-CoT w/ G.T. BBoxes in DocVQA and InfographicsVQA, illustrating the upper of our methods.
  • Both UV-CoT and UV-CoT perform better than Visual-CoT-7B across all zero-shot datasets. This demonstrates the strong generalization ability of our method in bounding box generation.
@@ 4.6 Insight of Evaluator Model
To gain a deeper understanding of the evaluator model, we conduct a series of variant experiments. The key findings, as presented in Tab. 5, are summarized as follows:
  • We compare UV-CoT with its self-evaluated variant, where the evaluator model is the same as the target model (initialed with LLaVA-1.5-7B). Although the self-evaluated version exhibits a performance decrease of 3.2% compared to the original UV-CoT, it still outperforms LLaVA-1.5-7B (+4.8% on average) across all evaluated datasets and achieves performance comparable to the larger OmniLMM-12B model (-0.2% on average). This demonstrates that UV-CoT maintains robust performance even under self-evaluation, highlighting its efficiency without requiring larger model scales.
  • For OmniLMM-12B and LLaVA-1.5-7B, we incorporate a CoT process using bounding boxes generated by UV-CoT. The CoT-enhanced versions significantly outperform their original counterparts, achieving average performance gains of 4.7% for LLaVA-1.5-7B and 7.4% for OmniLMM-12B across datasets. Remarkably, these models were not fine-tuned for the CoT process, indicating that the bounding box information alone substantially improves performance. This finding underscores that our evaluating process simplifies the task of generating complex spatial annotations, enabling MLLMs to focus on evaluating final answers.
@@ 4.7 Other Detailed Analyses
Visual Token Efficiency. Compared to standard answer generation in MLLMs, Visual CoT [28] and our UV-CoT reasoning double the number of visual tokens by processing additional local image regions. To evaluate token efficiency, i.e., performance under the same visual token budget, we resize the input image to different resolutions ($224^{2},336^{2}$ and $448^{2}$) and report the average score of different models in Fig. 4. (1) MLLMs with image-level reasoning (Visual CoT and our UV-CoT) demonstrate better token efficiency than the standard answer generation pipeline. Specifically, they achieve higher performance with 512 visual tokens than the standard pipeline does with 1024 tokens. (2) Our UV-CoT consistently outperforms the Visual CoT pipeline across all scales, achieving higher average scores due to it well-designed unsupervised learning pipeline. This highlights the token efficiency of our method.
Visualization. This section visualizes preference data and selected cases from our UV-CoT inference. As shown in Fig. 5, given different local regions and their corresponding answers, the evaluator MLLM assigns reasonable scores, validating the effectiveness of our automatic generation and labeling process. In Fig. 6, we present selected reasoning cases with model-generated bounding boxes overlaid. The precision of bounding box detection and the depth of understanding play a crucial role in determining the quality of generated answers.
@@ 5 Conclusion
In this work, we propose UV-CoT, a novel framework that enables image-level CoT reasoning in MLLMs via preference optimization techniques. Unlike previous methods that rely on SFT needing large amounts of labeled data, our approach leverages an unsupervised learning method to refine the model’s ability with image-level CoT using model-generated preference data (which are rough but useful). We address key challenges in preference data generation and effective optimization, ensuring a more adaptive and interpretable reasoning process. Extensive experiments demonstrate that UV-CoT achieves state-of-the-art performance, significantly improving visual comprehension in MLLMs on four reasoning datasets. Our findings from extensive experiments highlight the potential of preference learning as a scalable alternative to traditional SFT, enabling more robust and data-efficient multimodal reasoning.

### v2

@@ 4.1 Setup
4.2 Comparison with State-of-the-Art Methods
@@
Explicitly incorporating visual cues proves beneficial for multimodal reasoning. For instance, MiniCPM-o-8B, which uses rule-based cropping to focus on salient regions, outperforms LLaVA-1.5-7B by an average of 3.5%. However, its reliance on heuristics limits adaptability. In contrast, image-level CoT models like Visual-CoT and our UV-CoT achieve greater performance gains, even surpassing the larger LLaVA-1.5-13B, by leveraging MLLMs to adaptively generate key visual regions. This highlights the superior effectiveness of image-level CoT reasoning.
UV-CoT fundamentally differs from distillation/pseudo-labeling. Unlike distillation, where student performance is typically bounded by the teacher, UV-CoT outperforms its evaluator OmniLMM-12B by 5.1% on average. This suggests that UV-CoT goes beyond simply mimicking a larger model. Direct generation of accurate bounding boxes remains a challenge for MLLMs due to the need for precise spatial localization. Instead, UV-CoT reformulates the task as ranking—an inherently simpler and more tractable problem—which leads to better performance.
@@
UV-CoT outperforms the supervised Visual-CoT. Despite using significantly less data (249K unlabeled vs. 376K labeled), UV-CoT outperforms Visual-CoT-7B on TextVQA (+1.3%) and VSR (+1.6%). Moreover, UV-CoT(10%) surpasses Visual-CoT-7B by 2.1% on average, with notable gains on TextVQA (+2.5%), GQA (+2.2%), and VSR (+2.1%) and achieves comparable or better performance on the remaining datasets. This validates the effectiveness of our high-quality preference data generation and enhanced preference optimization method.
4.3 Zero-Shot Generalization
We evaluate the zero-shot performance of UV-CoT on the test sets of SROIE, DUDE, Visual7W, and V Bench [41], without any training exposure to these datasets. V Bench is a high-resolution benchmark (avg. size: $2246\times 1582$) covering diverse tasks; we focus on three representative ones: Attributes (object attribute recognition), GPT4V-Hard (complex visual reasoning), and OCR. Additionally, we train a variant of our model, UV-CoT, using preference data generated from the training splits of SROIE, DUDE, and Visual7W. The results are reported in Tab. 2 and Tab. 3, leading to the following observations:
@@ 4.3 Zero-Shot Generalization
UV-CoT exhibits stronger zero-shot performance. Supervised CoT learning relies on labeled data and often overfits to specific annotation distributions, limiting its generalization to unseen tasks. In contrast, UV-CoT uses preference optimization based on relative comparisons, avoiding reliance on absolute labels and enhancing generalization. As a result, UV-CoT outperforms all baselines across zero-shot datasets (+2.5% on average), with notable gains on DUDE and Visual7W (both +3.5%). Furthermore, UV-CoT achieves greater gains (5.1% on average), showing that our model can effectively learn the image-level CoT process without the need for costly human annotation.
UV-CoT excels in high-resolution image reasoning. Image-level CoT methods—UV-CoT (+19.8%) and Visual-CoT-7B (+14.3%)—significantly outperform non-CoT baselines, with over 50% gains on OCR tasks. UV-CoT further surpasses Visual-CoT-7B by 5.5% on average, achieving the best performance across all tasks. This large performance gap underscores the advantage of unsupervised image-level CoT for high-resolution visual reasoning.
Figure 3: (a&b) Bounding box evaluation on (a) training datasets and (b) zero-shot datasets. Our UV-CoT performs better than Visual-CoT. (c) Model performance under varying visual token sizes. Our UV-CoT demonstrates better token efficiency.
Model DocVQA TextVQA InfographicsVQA Flickr30k GQA VSR Average
OmniLMM-12B 0.254 0.578 0.172 0.621 0.509 0.523 0.443
IF: OmniLMM-12B + CoT 0.305 0.722 0.217 0.675 0.580 0.595 0.516
LLaVA-1.5-7B 0.198 0.507 0.131 0.539 0.480 0.504 0.393
IF: LLaVA-1.5-7B + CoT 0.245 0.577 0.149 0.606 0.530 0.533 0.440
Tr: UV-CoT (Evaluator: self-evaluated) 0.242 0.609 0.153 0.598 0.517 0.526 0.441
Tr: UV-CoT (Evaluator: OmniLMM-12B) 0.265 0.686 0.173 0.632 0.536 0.548 0.473
Table 5: Analysis of evaluator model. ‘+ CoT’ refers to inference assisted by bounding boxes generated by UV-CoT. Self-evaluated denotes using target model as the evaluator during training. IF and Tr indicate experiments conducted on the inference and training process.
@@ 4.4 Ablation Studies
In Tab. 4, we present the ablations on key components.
Image-level CoT: We design two variants to evaluate the impact of image-level CoT: (1) ‘w/o UV-CoT’ removes intermediate reasoning and directly outputs answers; (2) ‘UV-CoT w/ G.T. BBox’ replaces predicted regions with ground truth bounding boxes to assess localization quality. Results show that removing CoT leads to a significant drop (–7.7% on average), confirming its necessity. On Flickr30k, UV-CoT matches the G.T. variant, suggesting accurate region selection. However, in DocVQA and InfographicsVQA, using G.T. boxes yields better performance, revealing the difficulty of precise localization. This highlights the potential of our method and suggests future work for improving region accuracy for complex tasks.
Score-DPO (sDPO): ‘w/ naive DPO’ applies standard DPO loss and shows degraded performance across all datasets (1.9% on average), especially on DocVQA (2.5%) and Flickr30k (2.6%), revealing its limitations for CoT learning. In contrast, our sDPO incorporates preference scores to better differentiate choices, yielding consistent gains.
Iterative learning: ‘w/o iterative learning’ generates all preference pairs in a single pass and trains only once, leading to a significant performance drop (3.5% on average). This underscores the importance of iterative learning in continuously aligning the preference data distribution with the model’s evolving policy throughout the training process.
Response evaluation: ‘w/o $\gamma$ model’ sets $\gamma=0$ during response evaluation stage, ignoring the next response when generating preference scores. We observe a significant performance drop (8.8% on average), particularly on TextVQA (17.2%), highlighting the difficulty MLLMs face in directly evaluating bounding boxes. This underscores the necessity of incorporating next response in our evaluation method.
Figure 4: Visualization of preference data generated by Algorithm 1. Preferred BBoxes are in red. Dis-preferred BBoxes are in blue.
@@
4.5 Bounding Box Evaluation
We compare the quality of bounding boxes learned from supervised and unsupervised strategies using GPT-4o as a scorer, on both training datasets (Fig. 3 (a)) and zero-shot datasets (Fig. 3 (b)). Our main observations are:
(1) Our UV-CoT outperforms Visual-CoT-7B, achieving higher scores in five of six datasets, supporting its superior performance in generating helpful bounding box.
(2) The bounding box quality is closely related to the final performance. Our model exhibits lower scores (below 0.210) for bounding box generation in DocVQA and InfographicsVQA, correlating with its reduced final scores in these datasets (below 0.290 in Tab. 1). It underscores the validity of evaluating bounding box quality through its impact on subsequent answers.
(3) Both UV-CoT and UV-CoT outperform Visual-CoT-7B across all zero-shot datasets, which illustrates the strong generalization of our method in bounding box generation.
@@ 4.6 Insight of Evaluator Model
We further perform studies to better understand the role of evaluator model. Key findings from Tab. 5 are:
(1) We compare UV-CoT with its self-evaluated variant, where the evaluator model is the same as the target model (initialed with LLaVA-1.5-7B). Although the self-evaluated version exhibits a performance decrease of 3.2% compared to the original UV-CoT, it still outperforms LLaVA-1.5-7B (+4.8% on average) across all evaluated datasets and achieves performance comparable to the larger OmniLMM-12B model (-0.2% on average). This demonstrates that UV-CoT maintains robust performance even under self-evaluation, highlighting its efficiency without requiring larger model scales.
(2) For OmniLMM-12B and LLaVA-1.5-7B, we incorporate a CoT process using bounding boxes generated by UV-CoT. The CoT-enhanced versions significantly outperform their original counterparts, achieving average performance gains of 4.7% for LLaVA-1.5-7B and 7.4% for OmniLMM-12B. Remarkably, these models were not fine-tuned for the CoT process, indicating that the bounding box information alone substantially improves performance. This finding underscores that our evaluating process simplifies the task of generating complex spatial annotations, enabling MLLMs to focus on evaluating final answers.
@@ 4.7 Other Detailed Analyses
Visual token efficiency. Compared to standard MLLM generation, image-level CoT doubles the number of visual tokens by processing additional local image regions. To evaluate token efficiency, i.e., performance under the same visual token budget, we resize the input image to different resolutions ($224^{2},336^{2}$ and $448^{2}$) and report the average score of different models in Fig. 3(c). Our key findings are:
(1) MLLMs with image-level reasoning (Visual-CoT-7B and our UV-CoT) demonstrate better token efficiency than the standard answer generation pipeline. E.g., they achieve higher performance with 512 visual tokens than the standard pipeline does with 1024 tokens.
(2) Our UV-CoT consistently outperforms Visual-CoT-7B across all scales, achieving higher average scores. This highlights the token efficiency of our method.
Visualization. Fig. 4 visualizes some preference data from our UV-CoT inference process. Given different local regions and their corresponding answers, the evaluator MLLM assigns reasonable scores, validating the effectiveness of our automatic generation and labeling process. In Fig. 5, we present reasoning cases with model-generated bounding boxes overlaid. The precision of bounding box detection and the depth of understanding play a crucial role in determining the quality of generated answers.
@@ 5 Conclusion
In this work, we propose UV-CoT, a framework that enables image-level CoT reasoning in MLLMs via preference optimization. Unlike previous methods that rely on SFT needing large amounts of labeled data, our approach leverages unsupervised learning to refine the model’s ability with image-level CoT using model-generated preference data (which are rough but useful). We address key challenges in preference data generation and effective optimization, ensuring a more adaptive and interpretable reasoning process. Extensive experiments demonstrate that UV-CoT achieves state-of-the-art performance, significantly improving visual comprehension in MLLMs on ten reasoning datasets. Our findings highlight the potential of preference learning as a scalable alternative to traditional SFT, enabling more robust and data-efficient multimodal reasoning.
Acknowledgements
This research is supported by the National Research Foundation, Singapore under its AI Singapore Programme (AISG Award No: AISG3-RP-2022-030). Thanks to Zichen Tian for his assistance with figure visualization throughout this work. The authors also thank the reviewers for their valuable comments and suggestions.